OpenStudy (anonymous):

Will fan and medal Use a graphing calculator or spreadsheet software to solve 4x + y + z = 1 8x – 4y – 7z = 2 5y + 9z = 3 using matrices A. (0.5, -3, 2) B. (1, 2, 3) C. (-3, 0.5, 2) D. (0.5, 3, 2)

OpenStudy (irishboy123):

these are 3 planes, yep: \(4x + y + z = 1 \\ 8x – 4y – 7z = 25 \\ 0 x + y + 9z = 3\) any idea what the "answer" means ? :p

OpenStudy (anonymous):

Don't you solve for x, y, and z?

OpenStudy (anonymous):

what grade is this?

OpenStudy (anonymous):

@pokamons

OpenStudy (anonymous):

It's algebra 2

OpenStudy (anonymous):

im in geometry but i can try

OpenStudy (anonymous):

oka

OpenStudy (anonymous):

im sorrry, idk this :(

OpenStudy (anonymous):

Hey it's all good. I don't know this either.

OpenStudy (anonymous):

But thanks for trying anyway

OpenStudy (irishboy123):

I think you are supposed to do something like this: +++++++++++++++++ import numpy as np a = np.array( [ [4,1,1], [8,-4,-7], [0,5,9] ] ) b = np.array([1,2,3]) x = np.linalg.solve(a, b) print x +++++++++++++++++ Connected to pydev debugger (build 141.1899) [ 0.5 -3. 2. ] it says: "graphing calculator or spreadsheet software "

OpenStudy (irishboy123):

it certainly assumes some knowledge if that is lacking, we can have a crack at solving just 2 of them: \(4x + y + z = 1; \\ 8x – 4y – 7z = 2\)

OpenStudy (irishboy123):

or these: \(8x – 4y – 7z = 2 \\ 5y + 9z = 3\)

OpenStudy (anonymous):

[0.5, -3, 2] is correct. I just guessed and got 100%

OpenStudy (irishboy123):

awesome, telepathic!!!

OpenStudy (anonymous):

Thank you for your help!